# Complexity Analysis of Normalizing Constant Estimation: from Jarzynski Equality to Annealed Importance Sampling and beyond

### v1

@@ 3 Problem Setting
We make two remarks regarding this criterion. First, similar to how taking the mean of i.i.d. estimates reduces variance, we show in Lem. 29 that the probability above can be boosted to any $1-\zeta$ using the median trick: obtaining $O\left(\log\frac{1}{\zeta}\right)$ i.i.d. estimates satisfying Eq. 5 and taking their median. Therefore, we focus on the task of obtaining a single estimate satisfying Eq. 5 hereafter. Second, Eq. 5 also allows us to quantify the complexity of estimating the free energy $F=-\log Z$, which is often of greater interest in statistical mechanics than the partition function $Z$. We show in App. G that estimating $Z$ with $O(\varepsilon)$ relative error and estimating $F$ with $O(\varepsilon)$ absolute error share the same complexity up to constants. Further discussion of this guarantee, including a literature review and the comparison with bias and variance, is deferred to App. G.
Recall that the rationale behind annealing involves a gradual transition from $\pi_{0}$, a simple distribution that is easy to sample from and estimate the normalizing constant, to $\pi_{1}=\pi$, the more complicated target distribution. Throughout this paper, we define a curve of probability measures
@@
where $V_{1}=V$ is the potential of $\pi$, and the normalizing constant $Z_{1}=Z$ is what we need to estimate. We do not specify the exact form of this curve now, but only introduce the following mild regularity assumption on the curve, as assumed in classical textbooks such as Ambrosio et al. (2008, 2021), Santambrogio (2015):
@@
The potential $[0,1]\times\mathbb{R}^{d}\ni(\theta,x)\mapsto V_{\theta}(x)\in\mathbb{R}$ is jointly $C^{1}$, and the curve $(\pi_{\theta})_{\theta\in[0,1]}$ is AC with finite action $\mathcal{A}:=\int_{0}^{1}\left|\dot{\pi}\right|_{\theta}^{2}\mathrm{d}\theta$.
@@
$V$ is $\beta$-smooth, and there exists $x_{*}$, with $\|x_{*}\|=:R\lesssim\frac{1}{\sqrt{\beta}}$ such that $\nabla V(x_{*})=0$. Moreover, let $m:=\sqrt{\operatorname{\mathbb{E}}_{\pi}\|\cdot\|^{2}}<{+\infty}$.
@@
Finding a global minimum of (possibly non-convex) $V$ is challenging, but it is always feasible to find some $x_{+}$ close to a stationary point $x_{*}$ using optimization algorithms (e.g., Allen-Zhu and Li (2018)) within negligible cost compared with the complexity for normalizing constant estimation. By considering the translated distribution $\pi(\cdot-x_{+})$, we can assume the existence of a stationary point near 00. The assumption $R\lesssim\frac{1}{\sqrt{\beta}}$ is for technical purposes in our proof.
Equipped with this foundational setup, we now proceed to introduce the annealed LD and annealed LMC algorithms, and establish an analysis for JE and AIS.
@@ 4 Analysis of the Jarzynski Equality
To elucidate how annealing works in the task of normalizing constant estimation, we first consider annealed Langevin diffusion (ALD), which runs LD with a dynamically changing target distribution. We introduce a reparameterized curve $(\widetilde{\pi}_{t}=\pi_{\frac{t}{T}})_{t\in[0,T]}$ for some large $T$ to be determined later, and define the ALD as the following SDE:
@@
The following Jarzynski equality provides a connection between the work functional and the free energy difference, which naturally yields a method for normalizing constant estimation.
@@
Let $\mathbb{P}^{\to}$ be the path measure of Eq. 6, and define the work functional $W$ and the free energy difference $\Delta F$ as
$W(X):=\frac{1}{T}\int_{0}^{T}\partial_{\theta}V_{\theta}|_{\theta=\frac{t}{T}}(X_{t})\mathrm{d}t,\qquad\Delta F:=-\log\frac{Z_{1}}{Z_{0}}.$
Then we have the following relation:
$\operatorname{\mathbb{E}}_{\mathbb{P}^{\to}}\mathrm{e}^{-W}=\mathrm{e}^{-\Delta F}.$
Below, we sketch the proof from Vargas et al. (2024, Prop. 3.3), which offers a crucial aspect for our analysis. The complete proof is detailed in Sec. C.1.
@@
Leveraging Girsanov theorem (Lem. 1) and Itô’s formula, one can establish the following identity of the RN derivative, known as the Crooks fluctuation theorem (Crooks, 1998, 1999):
@@
Assume for the moment that (i) $Z_{0}$ is known, (ii) we can exactly simulate Eq. 6, and (iii) we can calculate the work functional $W(X)$ given any continuous trajectory $X$. According to Thm. 5, $\widehat{Z}:=Z_{0}\mathrm{e}^{-W(X)}$ with $X\sim\mathbb{P}^{\to}$ is an unbiased estimator of $Z=Z_{0}\mathrm{e}^{-\Delta F}$. We establish an upper bound on the time $T$ required to run the ALD in order to satisfy the accuracy criterion Eq. 5 in the following theorem, whose proof is detailed in Sec. C.2.
@@
Under Assump. 1, it suffices to choose $T=\frac{32\mathcal{A}}{\varepsilon^{2}}$ to obtain $\operatorname{Pr}\left(\left|\frac{\widehat{Z}}{Z}-1\right|\leq\varepsilon\right)\geq\frac{3}{4}$.
@@
To illustrate the proof idea of Thm. 6, note that while the ALD (Eq. 6) targets the distribution $\widetilde{\pi}_{t}$ at time $t$, there is always a lag between $\widetilde{\pi}_{t}$ and the actual law of $X_{t}$. Similarly, the backward SDE (Eq. 7) can also be seen as a time-reversed ALD which targets $\widetilde{\pi}_{t}$ at time $t$, and the same lag exists. This lag turns out to be the source of the error in the estimator $\widehat{Z}$.
@@
In practice, to alleviate the issue of high variance in estimating free energy differences, Vaikuntanathan and Jarzynski (2008) proposed adding a compensatory drift term $v_{t}(X_{t})$ to the ALD (Eq. 6). Ideally, the optimal choice would eliminate the lag entirely, ensuring $X_{t}\sim\widetilde{\pi}_{t}$ for all $t\in[0,T]$. Inspired by this, we compare the path measure of ALD $\mathbb{P}^{\to}$ to the SDE having the perfect compensatory drift term, whose path measure $\mathbb{P}$ has marginal distribution $\widetilde{\pi}_{t}$ at time $t$. To make possible the perfect match, $v_{t}$ must satisfy the Fokker-Planck equation. The Girsanov theorem (Lem. 1) enables the computation of $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\to})$ and $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\leftarrow})$, which are related to $\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}^{2}$. Finally, among all admissible drift terms $v_{t}$, Lem. 2 suggests the optimal choice of $v^{*}_{t}=\lim_{\delta\to 0}\frac{T_{\widetilde{\pi}_{t}\to\widetilde{\pi}_{t+\delta}}-\operatorname{id}}{\delta}$ to minimize this norm, thereby leading to the metric derivative $|\dot{\widetilde{\pi}}|_{t}$ and the action $\mathcal{A}$. Through this approach, we derive a bound not explicitly relying on isoperimetric assumptions.
@@
A similar connection between free energy and action integral was discovered in stochastic thermodynamics (Sekimoto, 2010, Seifert, 2012), one paradigm for non-equilibrium thermodynamics. By the second law of thermodynamics, the averaged dissipated work, defined as the averaged work minus the free energy difference, i.e., $\mathcal{W}_{\mathrm{diss}}:=\mathcal{W}-\Delta F:=\operatorname{\mathbb{E}}_{\mathbb{P}^{\to}}W-\Delta F$, is non-negative. When the underlying process is modeled by an overdamped LD, $\mathcal{W}_{\mathrm{diss}}$ can be quantified by an action integral divided by the length of the process (Aurell et al., 2011, Chen et al., 2020). This follows from the observation that $\mathcal{W}_{\mathrm{diss}}=\operatorname{KL}(\mathbb{P}^{\to}\|\mathbb{P}^{\leftarrow})$ and then a similar argument to that above. This connection provides a finer description of the second law of thermodynamics (Aurell et al., 2012) over a finite time horizon. Finally, we also observe that our bound aligns with the $O\left(\frac{1}{T}\right)$ decay rate of the variance of the work in Mazonka and Jarzynski (1999) (see also Lelièvre et al. (2010, Chap. 4.1.4)), computed when the curve consists of Gaussian distributions with linearly varying means.
@@ 5 Analysis of the Annealed Importance Sampling
Suppose we have probability distributions $\pi_{\ell}=\frac{1}{Z_{\ell}}f_{\ell}$, $\ell\in\left[\!\left[0,M\right]\!\right]$ and transition kernels $F_{\ell}(x,\cdot)$, $\ell\in\left[\!\left[1,M\right]\!\right]$, and assume that each $\pi_{\ell}$ is an invariant distribution of $F_{\ell}$, $\ell\in\left[\!\left[1,M\right]\!\right]$. Define the path measure
@@
It can be easily demonstrated, as in Eq. 8, that $\log\frac{\mathrm{d}\mathbb{P}^{\to}}{\mathrm{d}\mathbb{P}^{\leftarrow}}(x_{0:M})=W(x_{0:M})-\Delta F$. Consequently, the identity $\operatorname{\mathbb{E}}_{\mathbb{P}^{\to}}{\frac{\mathrm{d}\mathbb{P}^{\leftarrow}}{\mathrm{d}\mathbb{P}^{\to}}}=1$ implies the desired equality. ∎
To study non-asymptotic complexity guarantees, we focus on a widely used curve in theoretical analysis (Brosse et al., 2018, Ge et al., 2020), which we refer to as the geometric interpolation [^2] Eq. 11 differs slightly from a widely used curve in applications (Gelman and Meng, 1998, Neal, 2001): $\pi_{\theta}\propto\pi^{1-\lambda(\theta)}\phi^{\lambda(\theta)}$, where $\phi$ is a prior distribution (typically Gaussian). We refer to both as geometric interpolation.:
@@
where $\lambda(\cdot)$ is a decreasing function with $\lambda(0)=2\beta$ and $\lambda(1)=0$, referred to as the annealing schedule. With this choice of $\lambda(0)$, by Assump. 2, the potential of $\pi_{0}$ is $\beta$-strongly-convex and $3\beta$-smooth, making sampling and normalizing constant estimation relatively easy. To estimate $Z_{0}$, we use the TI algorithm from Ge et al. (2020), which requires $\widetilde{O}\left(\frac{d^{\frac{3}{2}}}{\varepsilon^{2}}\right)$ gradient oracle calls. In a nutshell, TI is an equilibrium method that constructs a series of intermediate distributions and estimates adjacent normalizing constant ratios via expectation under these intermediate distributions, realized through MCMC sampling from each intermediate distribution. As TI is peripheral to our primary focus, we defer its full description and complexity analysis to Secs. H.1 and 19.
Given the curve Eq. 11, we introduce discrete time points $10=\theta_{0}<\theta_{1}<...<\theta_{M}=$ 1 to be specified later, and adopt the framework outlined in Thm. 7 by setting $\pi_{\ell}=\frac{1}{Z_{\ell}}f_{\ell}$ to correspond to $\pi_{\theta_{\ell}}=\frac{1}{Z_{\theta_{\ell}}}f_{\theta_{\ell}}$, albeit with a slight abuse of notation. To estimate the normalizing constant, we need to sample from the forward path measure $\mathbb{P}^{\to}$, and calculate the work function along the trajectory. Since $\pi_{\theta_{\ell}}$ must be an invariant distribution of the transition kernel $F_{\ell}$ in $\mathbb{P}^{\to}$, we define $F_{\ell}$ via running LD targeting $\pi_{\theta_{\ell}}$ for a short time $T_{\ell}$, i.e., $F_{\ell}(x,\cdot)$ is given by the law of $X_{T_{\ell}}$ in the following SDE initialized at $X_{0}=x$:
@@
In this setting, AIS can be interpreted as a discretized version of JE (Lelièvre et al., 2010, Remark 4.5). However, in practice, exact samples from $\pi_{0}$ are often unavailable, and the simulation of LD cannot be performed perfectly. To capture these practical considerations, we define the following sampling path measure:
^$\widehat{\mathbb{P}}^{\to}(x_{0:M})=\widehat{\pi}_{0}(x_{0})\prod_{\ell=1}^{M}\widehat{F}_{\ell}(x_{\ell-1},x_{\ell}),$ (13)
@@
We refer readers to Algorithm 1 for a summary of the detailed implementation of our proposed AIS algorithm, including the TI procedure and the update rules in Eq. 14. The following theorem delineates the oracle complexity of the algorithm required to obtain an estimate $\widehat{Z}$ meeting the desired accuracy criterion (Eq. 5), whose detailed proof can be located in App. D.
@@
Let $\widehat{Z}$ be the AIS estimator described as in Algorithm 1, i.e., $\widehat{Z}:=\widehat{Z}_{0}\mathrm{e}^{-W(x_{0:M})}$ where $\widehat{Z}_{0}$ is estimated by TI and $x_{0:M}\sim\widehat{\mathbb{P}}^{\to}$. Under Assumps. 2 and 1, consider the annealing schedule $\lambda(\theta)=2\beta(1-\theta)^{r}$ for some $1\leq r\lesssim 1$. Use $\mathcal{A}_{r}$ to denote the action of $(\pi_{\theta})_{\theta\in[0,1]}$ to emphasize the dependence on $r$. Then, the oracle complexity for obtaining an estimate $\widehat{Z}$ that satisfies the criterion $\operatorname{Pr}\left(\left|\frac{\widehat{Z}}{Z}-1\right|\leq\varepsilon\right)\geq\frac{3}{4}$ is
$\widetilde{O}\left(\frac{d^{\frac{3}{2}}}{\varepsilon^{2}}\vee\frac{m\beta\mathcal{A}_{r}^{\frac{1}{2}}}{\varepsilon^{2}}\vee\frac{d\beta^{2}\mathcal{A}_{r}^{2}}{\varepsilon^{4}}\right).$ (15)
We present a high-level proof sketch using Fig. 1. The continuous dynamics, comprising the forward path $\mathbb{P}^{\to}$, the backward path $\mathbb{P}^{\leftarrow}$, and the reference path $\mathbb{P}$, are depicted as three black curves. To address discretization error, the $\ell$-th red (purple) arrow proceeding from left to right represents the transition kernel $\widehat{F}_{\ell}$ ($B_{\ell}$), whose composition forms $\widehat{\mathbb{P}}^{\to}$ ($\mathbb{P}^{\leftarrow}$).
@@
Figure 1: Illustration of the proof idea for Thm. 8.
  1. 1.
@@
    Analogously to the analysis of JE (Thm. 6), define the reference path measure $\mathbb{P}$ with transition kernels $F^{*}_{\ell}$ such that $x_{\ell}\sim\pi_{\theta_{\ell}}$. Given the sampling path measure $\widehat{\mathbb{P}}^{\to}$, define $\overline{\mathbb{P}}^{\to}$ as the version of $\widehat{\mathbb{P}}^{\to}$ without the initialization error, i.e., by replacing $\widehat{\pi}_{0}$ with $\pi_{0}$ in Eq. 13.
  2. 2.

### v2

@@ 2 Preliminaries and Problem Setting
We make two remarks regarding ˜1. First, similar to how taking the mean of i.i.d. estimates reduces variance, we show in Lem.˜11 that the probability above can be boosted to $1-\zeta$, $\forall\zeta\in\left(0,\frac{1}{4}\right)$ using the median trick: obtaining $O\left(\log\frac{1}{\zeta}\right)$ i.i.d. estimates satisfying ˜1 and taking their median. Therefore, we focus on the task of obtaining a single estimate satisfying ˜1 hereafter. Second, ˜1 also allows us to quantify the complexity of estimating the free energy $F=-\log Z$, which is often of greater interest in statistical mechanics than the partition function $Z$. We show in App.˜G that estimating $Z$ with $O(\varepsilon)$ relative error and estimating $F$ with $O(\varepsilon)$ absolute error share the same complexity up to constants. Further discussion of this guarantee, including a literature review and the comparison with bias and variance, is deferred to App.˜G.
A straightforward method for estimating $Z$ is through importance sampling, i.e., $Z=\operatorname{\mathbb{E}}_{\pi_{0}}\frac{\widehat{\pi}}{\pi_{0}}$ for some tractable proposal distribution $\pi_{0}$, yet its variance can be large due to the mismatch between $\pi_{0}$ and $\pi$. The rationale behind annealing involves a gradual transition from $\pi_{0}$ to $\pi_{1}=\pi$. Throughout this paper, we consider a curve of probability measures denoted as
@@
where $V_{1}=V$ is the potential of $\pi$, and $Z_{1}=Z$ is what we need to estimate. We do not specify the exact form of this curve now, but only introduce the following mild regularity assumption on the curve, as assumed in classical textbooks such as Ambrosio et al. (2008, 2021), Santambrogio (2015):
@@
The potential $0,1]\times\mathbb{R}^{d}\ni(\theta,x)\mapsto V_{\theta}(x)\in\mathbb{R}$ is jointly $C^{1}$, and the curve $(\pi_{\theta})_{\theta\in[0,1]}$ is absolute continuous with finite action $\mathcal{A}:=\int_{0}^{1}\left|\dot{\pi}\right|_{\theta}^{2}\mathrm{d}\theta$.
Here, $|\dot{\pi}|_{\theta}:=\lim_{\delta\to 0}\frac{W_{2}(\pi_{\theta+\delta},\pi_{\theta})}{|\delta|}$ is the Wasserstein-2 ($\text{W}_{\text{2}}$) metric derivative of the curve $(\pi_{\theta})_{\theta\in[0,1]}$ at $\theta$, which measures the “speed” of the curve in the space of probability distributions, and absolute continuity means the above limit exists and is finite for all $\theta\in[0,1]$. A curve having a finite action is a weaker condition than requiring each $\pi_{\theta}$ to satisfy isoperimetric inequalities (e.g., Poincaré or log-Sobolev). We refer readers to Sec.˜A.2 for details of optimal transport (OT), and only remark the following lemma below, which connects the metric derivative with the continuity equation and serves a key tool in our analysis:
Lemma 1 (Informal version of Lem.˜5).
For any vector field $v=(v_{t})_{t\in[a,b]}$ on $\mathbb{R}^{d}$ that generates an absolute continuous curve of probability measures $\rho=(\rho_{t})_{t\in[a,b]}$ (i.e., the continuity equation $\partial_{t}\rho_{t}+\nabla\cdot(\rho_{t}v_{t})=0$, $\forall t\in[a,b]$ holds), we have $|\dot{\rho}|_{t}\leq\|v_{t}\|_{L^{2}(\rho_{t})}$, $\forall t\in[a,b]$. Furthermore, there exists a unique vector field that generates $\rho$ and satisfies the equality for all $t\in[a,b]$.
@@
$V$ is $\beta$-smooth (i.e., $\nabla V$ is $\beta$-Lipschitz), $\nabla V(0)=0$, and $m:=\sqrt{\operatorname{\mathbb{E}}_{\pi}\|\cdot\|^{2}}<{+\infty}$.
@@
One can always find a stationary point $x_{*}$ of (possibly non-convex) $V$ using optimization methods within negligible cost compared with the complexity for estimating $Z$. By considering the translated distribution $\pi(\cdot-x_{*})$, we assume 0 is a stationary point without loss of generality.
Equipped with this fundamental setup, we now proceed to introduce the JE and AIS, and establish an analysis for their complexity.
@@ 3 Analysis of the Jarzynski Equality
To elucidate how annealing works in the task of normalizing constant estimation, we first consider annealed Langevin diffusion (ALD), which runs Langevin diffusion (LD) with a dynamically changing target distribution. We introduce a reparameterized curve $(\widetilde{\pi}_{t}=\pi_{t/T})_{t\in[0,T]}$ for some large $T$ to be determined later, and define the ALD as the following SDE:
@@
The following Jarzynski equality provides a connection between the work functional and the free energy difference, which naturally yields an estimator of normalizing constant.
@@
Let $\mathbb{P}^{\to}$ be the path measure of ˜2. Then the work functional $W$ and the free energy difference $\Delta F$ have the following relation:
$\operatorname{\mathbb{E}}_{\mathbb{P}^{\to}}\mathrm{e}^{-W}=\mathrm{e}^{-\Delta F},\quad\text{where}~~W(X):=\frac{1}{T}\int_{0}^{T}\partial_{\theta}V_{\theta}|_{\theta=\frac{t}{T}}(X_{t})\mathrm{d}t~~\text{and}~~\Delta F:=-\log\frac{Z_{1}}{Z_{0}}.$
Below, we sketch the proof from Vargas et al. (2024, Prop. 3.3), which offers a crucial aspect for our analysis: the forward and backward SDEs (see also Sec.˜A.1). The complete proof is in Sec.˜C.1.
@@
Leveraging the Girsanov theorem (Lem.˜2) and Itô’s formula, one can establish the following identity of the RN derivative, known as the Crooks fluctuation theorem (Crooks, 1998, 1999):
@@
Under the ideal setting where (i) $Z_{0}$ is known, (ii) the ALD in ˜2 can be simulated exactly, and (iii) the work functional $W(X)$ can be computed precisely, Thm.˜1 provides an unbiased estimator $\widehat{Z}:=Z_{0}\mathrm{e}^{-W(X)}$ for $Z=Z_{0}\mathrm{e}^{-\Delta F}$. Despite its dominant use (Chipot and Pohorille, 2007, Lelièvre et al., 2010), the statistical efficiency of this estimator is not well understood. While it is known that the variance of $\widehat{Z}$ can be large, non-asymptotic analyses quantifying its efficiency is lacking. We address this gap by establishing an upper bound on the time $T$ required for the ALD to satisfy the accuracy criterion ˜1 in the following theorem, whose proof is detailed in Sec.˜C.2.
@@
Under Assump.˜1, it suffices to choose $T=\frac{32\mathcal{A}}{\varepsilon^{2}}$ to obtain $\operatorname{Pr}\left(\left|\frac{\widehat{Z}}{Z}-1\right|\leq\varepsilon\right)\geq\frac{3}{4}$.
We first observe that our bound aligns with the decay rate of the variance of the work in Mazonka and Jarzynski (1999) (see also Lelièvre et al. (2010, Chap. 4.1.4)), which considered a special case $\pi_{\theta}=\operatorname{\mathcal{N}}\left(\theta L,\frac{1}{K}\right)$. They showed that $W\sim\operatorname{\mathcal{N}}\left(B_{T},2B_{T}\right)$ with $B_{T}=\frac{L^{2}}{T}\left(1-\frac{1-\mathrm{e}^{-KT}}{KT}\right)$, and hence the normalized variance $\operatorname{Var}_{\mathbb{P}^{\to}}\frac{\widehat{Z}}{Z}=\mathrm{e}^{2B_{T}}-1$ is asymptotically $O\left(\frac{1}{T}\right)$ as $T\to\infty$. Our bound, under a different criterion ˜1, is $O\left(\frac{1}{T}\right)$ for all $T>0$.
@@
To illustrate the proof idea of Thm.˜2, note that while the ALD ˜2 targets the distribution $\widetilde{\pi}_{t}$ at time $t$, there is always a lag between $\widetilde{\pi}_{t}$ and the actual law of $X_{t}$. Similarly, the backward SDE ˜3 can also be seen as a time-reversed ALD starting from $T$ that targets $\widetilde{\pi}_{t}$ at time $t$, and the same lag exists. This lag turns out to be the source of the error in the estimator $\widehat{Z}$.
@@
In practice, to alleviate the issue of high variance in estimating free energy differences, Vaikuntanathan and Jarzynski (2008) proposed adding a compensatory drift term $v_{t}(X_{t})$ to the ALD ˜2. Ideally, the optimal choice would eliminate the lag entirely, ensuring $X_{t}\sim\widetilde{\pi}_{t}$ for all $t\in[0,T]$. Inspired by this, we compare the path measure of ALD $\mathbb{P}^{\to}$ to the SDE having the perfect compensatory drift term, whose path measure $\mathbb{P}$ has marginal distribution $\widetilde{\pi}_{t}$ at time $t$. To make possible the perfect match, it turns out that $v_{t}$ must satisfy the Fokker-Planck equation with $\widetilde{\pi}_{t}$. The Girsanov theorem (Lem.˜2) enables the computation of $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\to})$ and $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\leftarrow})$, which are related to $\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}^{2}$. Finally, among all admissible drift terms $v_{t}$, Lem.˜1 suggests an optimal choice of $v^{*}_{t}$ to minimize this norm, thereby leading to the metric derivative $|\dot{\widetilde{\pi}}|_{t}$ and the action $\mathcal{A}$. This way avoids the explicit dependence of isoperimetric assumptions in our bound.
@@
A similar connection between free energy and action integral was discovered in stochastic thermodynamics (Sekimoto, 2010, Seifert, 2012), one paradigm for non-equilibrium thermodynamics. By the second law of thermodynamics, the averaged dissipated work, defined as the averaged work minus the free energy difference, i.e., $\mathcal{W}_{\mathrm{diss}}:=\mathcal{W}-\Delta F:=\operatorname{\mathbb{E}}_{\mathbb{P}^{\to}}W-\Delta F$, is non-negative. When the underlying process is modeled by an overdamped LD, $\mathcal{W}_{\mathrm{diss}}$ can be quantified by an action integral divided by the time duration (Aurell et al., 2011, Chen et al., 2020). This follows from the observation that $\mathcal{W}_{\mathrm{diss}}=\operatorname{KL}(\mathbb{P}^{\to}\|\mathbb{P}^{\leftarrow})$ and then a similar argument to that above. This connection provides a finer description of the second law of thermodynamics (Aurell et al., 2012) over a finite time horizon.
Finally, we place Thm.˜2 within the broader theme of sampling v.s. normalizing constant estimation by comparing Thm.˜2 with the complexity of non-log-concave sampling. Guo et al. (2025) proved that under the same assumptions, the ALD ˜2 can draw a sample within $\varepsilon^{2}$-error in $\operatorname{KL}(\pi\|\cdot)$ with the same order of time $T\asymp\frac{\mathcal{A}}{\varepsilon^{2}}$. While the classical work Jerrum et al. (1986) proved the existence of a polynomial-time algorithm for sampling and a polynomial-time algorithm for estimating normalizing constant imply each other in the discrete settings, we establish a similar quantitative connection between the complexities of these two tasks in the continuous settings without log-concavity, opening a new avenue of research on understanding their relationship. Though reaching similar results, the proof strategies are different: Guo et al. (2025) is a direct application of Girsanov theorem between $\mathbb{P}^{\to}$ and $\mathbb{P}$, while Thm.˜2 involves more complicated backward SDE arguments.
@@ 4 Analysis of the Annealed Importance Sampling
Suppose we have probability distributions $\pi_{\ell}=\frac{f_{\ell}}{Z_{\ell}}$, $\ell\in\left[\!\left[0,M\right]\!\right]$ and transition kernels $F_{\ell}(x,\cdot)$, $\ell\in\left[\!\left[1,M\right]\!\right]$, and assume that each $\pi_{\ell}$ is an invariant distribution of $F_{\ell}$, $\ell\in\left[\!\left[1,M\right]\!\right]$. Define the path measure
@@
It can be easily demonstrated, as in ˜4, that $\log\frac{\mathrm{d}\mathbb{P}^{\to}}{\mathrm{d}\mathbb{P}^{\leftarrow}}(x_{0:M})=W(x_{0:M})-\Delta F$. Consequently, the identity $\operatorname{\mathbb{E}}_{\mathbb{P}^{\to}}{\frac{\mathrm{d}\mathbb{P}^{\leftarrow}}{\mathrm{d}\mathbb{P}^{\to}}}=1$ implies the desired equality. ∎
While JE and AIS holds for general interpolations, for the study non-asymptotic complexity guarantees, we focus on a widely used curve in theoretical analysis (Brosse et al., 2018, Ge et al., 2020), which we refer to as the geometric interpolation:2227 differs slightly from a widely used curve in applications (Gelman and Meng, 1998, Neal, 2001): $\pi_{\theta}\propto\pi^{1-\lambda(\theta)}\phi^{\lambda(\theta)}$, where $\phi$ is a prior distribution (typically Gaussian). We refer to both as geometric interpolation.
@@
where $\lambda(\cdot)$ is a decreasing function with $\lambda(0)=2\beta$ and $\lambda(1)=0$, referred to as the annealing schedule. With this choice of $\lambda(0)$, by Assump.˜2, the potential of $\pi_{0}$ is $\beta$-strongly-convex and $3\beta$-smooth, making sampling and normalizing constant estimation relatively easy. To estimate $Z_{0}$, we use the TI algorithm from Ge et al. (2020), which requires $\widetilde{O}\left(\frac{d^{4/3}}{\varepsilon^{2}}\right)$ gradient oracle calls. In a nutshell, TI is an equilibrium method that constructs a series of intermediate distributions and estimates adjacent normalizing constant ratios via expectation under these intermediate distributions, realized through MCMC sampling from each intermediate distribution. As TI is peripheral to our primary focus, we defer its full description to Sec.˜H.1.
Given ˜7, we introduce time points $0=\theta_{0}<\theta_{1}<...<\theta_{M}=1$ to be specified later, and adopt the framework outlined in Thm.˜3 by setting $\pi_{\ell}=\frac{f_{\ell}}{Z_{\ell}}$ to correspond to $\pi_{\theta_{\ell}}=\frac{f_{\theta_{\ell}}}{Z_{\theta_{\ell}}}$, albeit with a slight abuse of notation. To estimate the normalizing constant, we need to sample from the forward path measure $\mathbb{P}^{\to}$ and compute the work function along the trajectory. Since $\pi_{\theta_{\ell}}$ must be an invariant distribution of the transition kernel $F_{\ell}$ in $\mathbb{P}^{\to}$, we define $F_{\ell}$ via running LD targeting $\pi_{\theta_{\ell}}$ for a short time $T_{\ell}$, i.e., $F_{\ell}(x,\cdot)$ is given by the law of $X_{T_{\ell}}$ in the following SDE initialized at $X_{0}=x$:
@@
In this setting, AIS can be interpreted as a discretization of JE (Lelièvre et al., 2010, Rmk. 4.5). However, in practice, exact samples from $\pi_{0}$ are often unavailable, and the simulation of LD cannot be performed perfectly. [^3] Recall that we need $F_{\ell}$ to have invariant distribution $\pi_{\ell}$ in Thm. 3. To capture these considerations, we define the following path measure:
$\widehat{\mathbb{P}}^{\to}(x_{0:M})=\widehat{\pi}_{0}(x_{0})\prod_{\ell=1}^{M}\widehat{F}_{\ell}(x_{\ell-1},x_{\ell}),$ (9)
@@
We refer readers to Algorithm˜1 in App.˜B for a summary of the detailed implementation of our proposed AIS algorithm, including the TI procedure and the update rules in ˜10. The following theorem delineates the oracle complexity of the algorithm required to obtain an estimate $\widehat{Z}$ meeting the desired accuracy criterion ˜1, whose detailed proof can be located in App.˜D.
@@
Let $\widehat{Z}$ be the AIS estimator described as in Algorithm˜1, i.e., $\widehat{Z}:=\widehat{Z}_{0}\mathrm{e}^{-W(x_{0:M})}$ where $\widehat{Z}_{0}$ is estimated by TI and $x_{0:M}\sim\widehat{\mathbb{P}}^{\to}$. Under Assumps.˜2 and 1, consider the annealing schedule $\lambda(\theta)=2\beta(1-\theta)^{r}$ for some $1\leq r\lesssim 1$. Use $\mathcal{A}_{r}$ to denote the action of $(\pi_{\theta})_{\theta\in[0,1]}$ to emphasize the dependence on $r$. Then, the oracle complexity for obtaining an estimate $\widehat{Z}$ that satisfies ˜1 is
$\widetilde{O}\left(\frac{d^{\frac{4}{3}}}{\varepsilon^{2}}\vee\frac{m\beta\mathcal{A}_{r}^{\frac{1}{2}}}{\varepsilon^{2}}\vee\frac{d\beta^{2}\mathcal{A}_{r}^{2}}{\varepsilon^{4}}\right).$ (11)
@@
Figure 1: Illustration of the proof idea for Thm.˜4.
We present a high-level proof sketch using Fig.˜1. The continuous dynamics, comprising the forward path $\mathbb{P}^{\to}$, the backward path $\mathbb{P}^{\leftarrow}$, and the reference path $\mathbb{P}$, are depicted as three black curves. To address discretization error, the $\ell$-th red (purple) arrow proceeding from left to right represents the transition kernel $\widehat{F}_{\ell}$ ($B_{\ell}$), whose composition forms $\widehat{\mathbb{P}}^{\to}$ ($\mathbb{P}^{\leftarrow}$).
@@
(I) Analogously to the analysis of JE (Thm.˜2), define the reference path measure $\mathbb{P}$ with transition kernels $F^{*}_{\ell}$ such that $x_{\ell}\sim\pi_{\theta_{\ell}}$. Given the sampling path measure $\widehat{\mathbb{P}}^{\to}$, define $\overline{\mathbb{P}}^{\to}$ as the version of $\widehat{\mathbb{P}}^{\to}$ without the initialization error, i.e., by replacing $\widehat{\pi}_{0}$ with $\pi_{0}$ in ˜9.